Let fn(x) =1/(1+n2x2) if 0 less than or equal to x less than or equal to 1,n=1,2,........ Prove that {fn} converges point wise ,but not uniformly on [0,1] Is term by term integration permissible?


1
Expert's answer
2021-03-10T10:03:21-0500

Here 

"lim_{n\rightarrow \infty} f_n(x)=0, \forall x"


Now "\dfrac{1}{1+n^2x^2}" attains maximum value "\dfrac{1}{2} \text{ at } x=\dfrac{1}{n}"


 tending to 0 as "n\rightarrow \infty."


Let us take the interval [0,1] containg 0.


Thus

 "M_n=\underset {x\in [a,b]}{Sup}|f_n(x)-f(x)|"


     "= \underset{x\in [a,b]}{Sup}|\dfrac{1}{1+n^2x^2}|"


     "=\dfrac{1}{2}" Which does not tends to zero as "n\rightarrow \infty."


 Hence The sequence is not uniformly convergent in interval [0,1].